OpenStudy (anonymous):

Solve the system by substitution. Check your solution. 4072-03-01-06-00_files/i0160000.jpg a. (15, 15) c. (13, 12) b. (10, 12) d. (7, 9)

OpenStudy (anonymous):

@i_love_bands26 @Lady.Liv1776 @ help

OpenStudy (igreen):

What's the system of equations?

OpenStudy (anonymous):

that

OpenStudy (anonymous):

just fanning you

OpenStudy (anonymous):

the m was a accident

OpenStudy (anonymous):

@Ziferot help

OpenStudy (igreen):

Okay. a - 1.2b = -3 0.2b + 0.6a = 12 a - 1.2b = -3 Add 1.2b to both sides: a = 1.2b - 3 Substitute 1.2b - 3 into the 2nd equation for 'a': 0.2b + 0.6a = 12 0.2b + 0.6(1.2b - 3) = 12 Distribute 0.6: 0.2b + 0.72b - 1.8 = 12 Add 0.2b + 0.72

OpenStudy (anonymous):

do i have to do that equation?

OpenStudy (anonymous):

brb

OpenStudy (igreen):

0.2b + 0.72b = ?

OpenStudy (igreen):

Do you know how to add that?

OpenStudy (anonymous):

just had to get a graphing calulator

OpenStudy (anonymous):

how to add letters

OpenStudy (igreen):

0.2b + 0.72b is the same as 0.2 + 0.72, but you just put the variable b next to it.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@Ziferot my page says you are on two questions

OpenStudy (igreen):

0.2 + 0.72 = ?

OpenStudy (anonymous):

how to put the variable in?

OpenStudy (igreen):

Just put it next to the number..like in 0.2b

OpenStudy (anonymous):

it says 0

OpenStudy (anonymous):

@Lady.Liv1776 do you know?

OpenStudy (igreen):

0.2 + 0.72 = ?

OpenStudy (anonymous):

i try that

OpenStudy (igreen):

What did you get?

OpenStudy (anonymous):

didn't work

OpenStudy (igreen):

Try again

OpenStudy (anonymous):

have to go

OpenStudy (anonymous):

be right back

OpenStudy (anonymous):

@igreen @Ggirl99 @i_love_bands26 @Lady.Liv1776 @Ziferot @undeadknight26 I AM BACK

OpenStudy (anonymous):

@trinity123 HELP

OpenStudy (anonymous):

@iGreen HELP

OpenStudy (igreen):

0.2b + 0.72b = 0.92b

OpenStudy (igreen):

So we have: 0.92b - 1.8 = 12 Add 1.8 to both sides, 12 + 1.8 = ?

OpenStudy (anonymous):

@chasity.ambs HELP

OpenStudy (igreen):

What's 12 + 1.8?

OpenStudy (anonymous):

K give me a minute

OpenStudy (anonymous):

@daiglejeremy101 what is the question

OpenStudy (anonymous):

Solve the system by substitution. Check your solution. http://media.education2020.com/evresources/4072-03-01-06-00_files/i0160000.jpg A. (15, 15) c. (13, 12) b. (10, 12) d. (7, 9)

OpenStudy (anonymous):

A is the answer i am pretty sure @daiglejeremy101

OpenStudy (anonymous):

THANK YOU @chasity.ambs

OpenStudy (igreen):

@chasity.ambs Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

Your welcome @daiglejeremy101 You can put this solution on YOUR website! here we have to find out values of "a" and "b" by substitution method and verify the solution. a-1.2b = -3 ...............Equation-1 0.2b+ 0.6a= 12 ...............Equation-2 From Equation-1: a = (-3 + 1.2b )= (1.2b - 3)...Equation-3 Now substitute "a" from above (Equation-3) into Equation-2 to get following equation: 0.2b + 0.6(1.2b-3)=12 0.2b + 0.72b - 1.8 = 12 0.92b - 1.8 = 12 0.92b = 12 + 1.8 = 13.8 b = 13.8/0.92 = 15 Therefore b = 15 Substitute b=15 into Equation-3 to get "a" as below: a = 1.2b - 3 = 1.2*15 - 3 = 18 - 3 = 15 Hence values of : a=15 & b=15 Verification of Solution: -------------------------- Substitute values of a & b into Equation-1 to get -3 as below: a-1.2b = 15 - 1.2*15 = 15 - 18 = -3 --------> a-1.2b= -3 Now substitute values of a & b into Equation-2 to get 12 as below: 0.2b+ 0.6a = 0.2*15 + 0.6*15 = 3 + 9 = 12 -------> 0.2b+0.6a=12 thats it.... B.Shridhar is that better @iGreen if you can show them how to do with out pen an paper be my guest!!!!

OpenStudy (anonymous):

CANT IT IS E2020

OpenStudy (anonymous):

I post that to make @iGreen Happy sorry to confuse you @daiglejeremy101

OpenStudy (anonymous):

THAT IS OK

OpenStudy (igreen):

@chasity.ambs You copy & pasted that..and even if you didn't you weren't supposed to give the answer out..we need participation of the Asker.

OpenStudy (anonymous):

I showed him how i got the answer so get off my CASE!!! @iGreen

OpenStudy (anonymous):

next question

OpenStudy (anonymous):

Jeannie is 8 years older than her younger sister April. When added together, their ages equal 198. How old is April? A. 27 B. 35 C. 95 D. 103

OpenStudy (anonymous):

you know this bohotness girl

OpenStudy (anonymous):

@GhostBirdSeriesNerd help please

OpenStudy (anonymous):

http://www.jiskha.com/ use this web site it may be some type of help.... I cant tell you answer so use the web site @daiglejeremy101

OpenStudy (anonymous):

thank you so much :) :)

OpenStudy (anonymous):

Any time i can help

OpenStudy (anonymous):

Cynthia has a cup full of dimes and quarters that she would like to put in her savings account. When she takes the cup to the bank, the banker tells Cynthia that the 386 coins in the cup are worth $58.10. How many quarters does Cynthia have in her cup? A. 126 B. 130 C. 256 D. 386

OpenStudy (anonymous):

hi @Ggirl99

OpenStudy (anonymous):

hi :D

OpenStudy (anonymous):

http://openstudy.com/study#/updates/518347f3e4b09587cd3904d6 This mite help....@daiglejeremy101

OpenStudy (anonymous):

A system of two linear equations has an infinite number of solutions. How is this possible? A. It's not; the graphs of the two equations can intersect only once. B. The two equations are the same line. C. The equations are parallel lines. D. There was an error in solving the system.